Webster World has sales of $12,900, costs of $5,800, depreciation expense of $1,100, and interest expense of $700. What is the operating cash flow if the tax rate is 32 percent? O $4.704 O $5,404 O $5,749 O $7.036
Answer #1
a sales 12900
b Cost 5800
c Depreciation 1100
d interest 700
e=a-b-c-d Earning before tax 5300
f=e*32% Tax @ 32% 1696
g=e-f Profit after tax 3604
h=g+d+c Operating cash flow = (profit after tax + Depreciation + interest) 5404
Correct answer is option : 5404